# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v2

@@ Appendix B Limitations
Appendix F Experimental Details
F.1 Simulation Details
@@
F.2 Experimental Details on LLMs
@@
Generating such counterfactual pairs, however, presents significant challenges. First, the inherent complexity and nuances of natural language make it difficult to create pairs that differ in precisely the intended contexts while leaving other aspects unchanged. Second, as highlighted in prior works (Park et al., 2023; Jiang et al., 2024), constructing such counterfactual sentences is a highly non-trivial task, even for human annotators, due to the intricacies of semantics and the need for precise control over contextual variations.
@@
To overcome these challenges, we adopt an approach based on existing resources. Specifically, we utilize the 27 counterfactual pairs introduced in (Park et al., 2023), which provide a structured and well-curated set of counterfactual pairs. These concepts encompass a wide range of semantic and morphological transformations, as detailed in Table 1. By leveraging this established dataset, we ensure consistency with previous research while facilitating a robust and meaningful evaluation of Corollary 4.2.
We first use these 27 counterfactual pairs to generate a submatrix of the linear transformation matrix $\mathbf{A}$. As stated in Corollary 4.1, the differences between the counterfactual pairs correspond to the columns of $\mathbf{A}$. As a result, we can create a submatrix $\mathbf{A}_{s}$ with size $27\times dim$by using the differences in the features of these 27 counterfactual pairs, where $dim$ corresponds to the feature dimension of the used LLM, such as 4096 for the Llama-27B model.
@@
To generate a submatrix of the weight matrix $\mathbf{W}$, we train a linear classifier using these 27 counterfactual pairs. Specifically, we use the features of the counterfactual pairs as input and the corresponding values of the latent variables as output. This allows us to obtain linear weights from the 27 counterfactual pairs, such as a sub-weight matrix $\mathbf{W}_{s}$. In our experiments, we employ the LogisticRegression classifier from the scikit-learn library.
@@
Note that, as mentioned in Corollary 4.1, the obtained submatrix $\mathbf{A}_{s}$ is scaled, meaning that $\mathbf{W}_{s}$ may also have a scaling factor. Therefore, we normalize both $\mathbf{A}_{s}$ and $\mathbf{W}_{s}$, before computing $\mathbf{A}_{s}\times\mathbf{W}_{s}$. The scaling indeterminacy of latent variables is an inherent property of the latent space. Without additional information to fix the scale within the latent space, the estimation can only be identified up to a scaling factor. However, in most cases, the exact scaling factor is not particularly significant. For example, concepts are often represented as directions rather than magnitudes.
@@
Table 1: Concept names, one example of the counterfactual pairs, and the number of used pairs, taken from (Park et al., 2023).
# Concept Example Count
1 verb $\Rightarrow$ 3pSg (accept, accepts) 32
2 verb $\Rightarrow$ Ving (add, adding) 31
3 verb $\Rightarrow$ Ved (accept, accepted) 47
4 Ving $\Rightarrow$ 3pSg (adding, adds) 27
5 Ving $\Rightarrow$ Ved (adding, added) 34
6 3pSg $\Rightarrow$ Ved (adds, added) 29
7 verb $\Rightarrow$ V + able (accept, acceptable) 6
8 verb $\Rightarrow$ V + er (begin, beginner) 14
9 verb $\Rightarrow$ V + tion (compile, compilation) 8
10 verb $\Rightarrow$ V + ment (agree, agreement) 11
11 adj $\Rightarrow$ un + adj (able, unable) 5
12 adj $\Rightarrow$ adj + ly (according, accordingly) 18
13 small $\Rightarrow$ big (brief, long) 20
14 thing $\Rightarrow$ color (ant, black) 21
15 thing $\Rightarrow$ part (bus, seats) 13
16 country $\Rightarrow$ capital (Austria, Vienna) 15
@@ 5 Ving $\Rightarrow$ Ved (adding, added) 34
18 male $\Rightarrow$ female (actor, actress) 11
19 lower $\Rightarrow$ upper (always, Always) 34
20 noun $\Rightarrow$ plural (album, albums) 63
21 adj $\Rightarrow$ comparative (bad, worse) 19
22 adj $\Rightarrow$ superlative (bad, worst) 9
23 frequent $\Rightarrow$ infrequent (bad, terrible) 32
24 English $\Rightarrow$ French (April, avril) 46
25 French $\Rightarrow$ German (ami, Freund) 35
26 French $\Rightarrow$ Spanish (annee, año) 35
27 German $\Rightarrow$ Spanish (Arbeit, trabajo) 22
Appendix G Further Discussion: Observations on LLMs and World Models
G.1 LLMs Mimic the Human World Model, Not the World Itself

### v3

@@ Appendix B Limitations
Appendix F Extension of Corollaries 4.2 and 4.3 for a Binary Concept
When considering pair that differ only in a binary concept of interest, i.e., $c^{i}$, both Corollaries 4.2 and 4.3 can be further refined, yielding the following results.
F.1 Extension of Corollary 4.2 for a Binary Concept
Corollary F.1 (Binary Concept Direction).
Suppose that Theorem 3.1 holds, and let $c^{i}$ be a binary concept variable, i.e., $c^{i}\in\{0,1\}$. Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ be a pair of inputs that differ only in the $i$-th binary concept $c^{i}$, with $c^{i}=0$ for $\mathbf{x}_{0}$ and $c^{i}=1$ for $\mathbf{x}_{1}$. Then, as $\epsilon_{\mathbf{x}}\to 0$ in Definition 4.1, the representation difference simplifies as:
$\mathbf{f}_{x}(\mathbf{x}_{1})-\mathbf{f}_{x}(\mathbf{x}_{0})\approx\tilde{\mathbf{A}}^{i}\left(\left[\log p(c^{i}\mid\mathbf{x}_{1})-\log p(c^{i}\mid\mathbf{x}_{0})\right]_{c^{i}}\right)$ $\approx\log p(c^{i}=0\mid\mathbf{x}_{1})\cdot\tilde{\mathbf{A}}^{i}\begin{bmatrix}1\\ -1\end{bmatrix},$ (65)
or, $\approx\log p(c^{i}=1\mid\mathbf{x}_{0})\cdot\tilde{\mathbf{A}}^{i}\begin{bmatrix}1\\ -1\end{bmatrix},$ (66)
where $\tilde{\mathbf{A}}^{i}=\mathbf{A}\mathbf{B}^{i}$, $\mathbf{B}^{i}$ is a binary lifting matrix that broadcasts each entry of $[\log p(c^{i}\mid\mathbf{x})]_{c^{i}}$ to the corresponding index in $[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})]_{\mathbf{c}_{i}}$. This shows that changes in a binary concept are encoded in a specific direction in the representation space defined by $\tilde{\mathbf{A}}^{i}$.
Proof.
Recall Eq. (36), the joint concept distribution conditioned on input $\mathbf{x}$ follows a peaked structure:
$p(\mathbf{c}\mid\mathbf{x})=\begin{cases}1-\epsilon_{\mathbf{x}},&\text{if }\mathbf{c}=\mathbf{c}^{*}_{\mathbf{x}},\\ \epsilon_{\mathbf{x}}\cdot r(\mathbf{c}),&\text{otherwise},\end{cases}$ (67)
where $\mathbf{c}^{*}_{\mathbf{x}}=(c^{i*},\mathbf{c}^{-i*})$ is the dominant concept configuration for $\mathbf{x}$, and $r(\mathbf{c})$ is a normalized residual distribution over all non-dominant $\mathbf{c}$.
Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ differ only in the $i$-th binary concept variable $c^{i}$, with:
$\mathbf{c}^{*}_{\mathbf{x}_{0}}=(0,\mathbf{c}^{-i*}),\quad\mathbf{c}^{*}_{\mathbf{x}_{1}}=(1,\mathbf{c}^{-i*}).$ (68)
Then the marginal probabilities for $c^{i}$ are:
For $\mathbf{x}_{1}$:
$p(c^{i}=1\mid\mathbf{x}_{1})=(1-\epsilon_{\mathbf{x}_{1}})+\epsilon_{\mathbf{x}_{1}}\cdot\sum_{\mathbf{c}^{-i}\neq\mathbf{c}^{-i*}}r(1,\mathbf{c}^{-i})=1-\epsilon_{\mathbf{x}_{1}}\cdot\alpha_{1},$ (69)
where $\alpha_{1}:=1-\sum_{\mathbf{c}^{-i}\neq\mathbf{c}^{-i*}}r(1,\mathbf{c}^{-i})$.
$p(c^{i}=0\mid\mathbf{x}_{1})=1-p(c^{i}=1\mid\mathbf{x}_{1})=\epsilon_{\mathbf{x}_{1}}\cdot\alpha_{1}.$ (70)
For $\mathbf{x}_{0}$:
$p(c^{i}=0\mid\mathbf{x}_{0})=(1-\epsilon_{\mathbf{x}_{0}})+\epsilon_{\mathbf{x}_{0}}\cdot\sum_{\mathbf{c}^{-i}\neq\mathbf{c}^{-i*}}r(0,\mathbf{c}^{-i})=1-\epsilon_{\mathbf{x}_{0}}\cdot\alpha_{0},$ (71)
where $\alpha_{0}:=1-\sum_{\mathbf{c}^{-i}\neq\mathbf{c}^{-i*}}r(0,\mathbf{c}^{-i})$.
$p(c^{i}=1\mid\mathbf{x}_{0})=1-p(c^{i}=1\mid\mathbf{x}_{0})=\epsilon_{\mathbf{x}_{0}}\cdot\alpha_{0}.$ (72)
Taking logarithmic, we have:
$\log p(c^{i}=1\mid\mathbf{x}_{1})=\log(1-\epsilon_{\mathbf{x}_{1}}\cdot\alpha_{1})\approx 0,\quad\text{as }\epsilon_{\mathbf{x}}\to 0$ (73)
$\log p(c^{i}=0\mid\mathbf{x}_{1})=\log(\epsilon_{\mathbf{x}_{1}}\cdot\alpha_{1}).$ (74)
$\log p(c^{i}=0\mid\mathbf{x}_{0})=\log(1-\epsilon_{\mathbf{x}_{0}}\cdot\alpha_{0})\approx 0,\quad\text{as }\epsilon_{\mathbf{x}}\to 0$ (75)
$\log p(c^{i}=1\mid\mathbf{x}_{0})=\log(\epsilon_{\mathbf{x}_{0}}\cdot\alpha_{0})\approx\log(\epsilon_{\mathbf{x}_{1}}\cdot\alpha_{1})=\log p(c^{i}=0\mid\mathbf{x}_{1}),\quad\text{as }\epsilon_{\mathbf{x}}\to 0.$ (76)
Then the vector difference:
$\left[\log p(c^{i}\mid\mathbf{x}_{1})-\log p(c^{i}\mid\mathbf{x}_{0})\right]_{c^{i}}$ $=\begin{bmatrix}\log p(c^{i}=0\mid\mathbf{x}_{1})-\log p(c^{i}=0\mid\mathbf{x}_{0})\\ \log p(c^{i}=1\mid\mathbf{x}_{1})-\log p(c^{i}=1\mid\mathbf{x}_{0})\end{bmatrix}$ (77)
$\approx\begin{bmatrix}\log p(c^{i}=0\mid\mathbf{x}_{1})\\ -\log p(c^{i}=1\mid\mathbf{x}_{0})\end{bmatrix}$ (78)
$\approx\log p(c^{i}=0\mid\mathbf{x}_{1})\begin{bmatrix}1\\ -1\end{bmatrix}$ (79)
Finally, the representation difference becomes:
$\mathbf{f}_{x}(\mathbf{x}_{1})-\mathbf{f}_{x}(\mathbf{x}_{0})\approx\tilde{\mathbf{A}}^{i}\left(\left[\log p(c^{i}\mid\mathbf{x}_{1})-\log p(c^{i}\mid\mathbf{x}_{0})\right]_{c^{i}}\right)\approx\log p(c^{i}=0\mid\mathbf{x}_{1})\cdot\tilde{\mathbf{A}}^{i}\begin{bmatrix}1\\ -1\end{bmatrix}.$ (80)
Here, note that: $\log p(c^{i}=0\mid\mathbf{x}_{1})\approx\log p(c^{i}=1\mid\mathbf{x}_{0})$ as shown in Eq. 76. ∎
F.2 Extension of Corollary 4.3 for a Binary Concept
Corollary F.2 (Binary Concept Classification).
Suppose that Theorem 3.1 holds, i.e., $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}+\mathbf{b}$. Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ be pair data that differ only in the $i$-th binary concept variable $c^{i}$, with labels $c^{i}$, where $c^{i}=0$ for $\mathbf{x}_{0}$ and $c^{i}=1$ for $\mathbf{x}_{1}$. Then when $\epsilon_{\mathbf{x}}\to 0$ in Def. 4.1, the corresponding representations $(\mathbf{f}(\mathbf{x}_{0}),\mathbf{f}(\mathbf{x}_{1}))$ are linearly separable with a weight vector $\mathbf{w}$ satisfying $\mathbf{w}\tilde{\mathbf{A}}^{i}\begin{bmatrix}1\\ -1\end{bmatrix}\approx s$, where $s$ accounts for softmax scaling. The corresponding logit is the unnormalized $p({c}^{i}=1\mid\mathbf{x})$.
Proof.
When $\epsilon_{\mathbf{x}}\to 0$, the posterior $p(\mathbf{c}\mid\mathbf{x})$ becomes sharply peaked at a unique mode $\mathbf{c}^{*}_{\mathbf{x}}$. This implies a near-independence of $c^{i}$ and $\mathbf{c}^{-i}$ given $\mathbf{x}$, allowing us to write:
$p(\mathbf{c}\mid\mathbf{x})\approx p(c^{i}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i}\mid\mathbf{x}).$ (81)
Taking logs and substituting into Theorem 3.1 (neglecting the bias term $\mathbf{b}$), we obtain:
$\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}\left[\log p(\mathbf{c}\mid\mathbf{x})\right]_{\mathbf{c}}\approx\mathbf{A}\left(\mathbf{B}^{i}[\log p(c^{i}\mid\mathbf{x})]_{c^{i}}+\mathbf{B}^{-i}[\log p(\mathbf{c}^{-i}\mid\mathbf{x})]_{\mathbf{c}^{-i}}\right),$ (82)
where $\mathbf{B}^{i}$ and $\mathbf{B}^{-i}$ denote the lifting operators that map the marginal log-probabilities of $c^{i}$ and $\mathbf{c}^{-i}$ into the joint log-probability vector space.
Define $\tilde{\mathbf{A}}^{i}:=\mathbf{A}\mathbf{B}^{i}$. Then:
$\mathbf{f}_{x}(\mathbf{x})\approx\tilde{\mathbf{A}}^{i}[\log p(c^{i}\mid\mathbf{x})]_{c^{i}}+\text{(independent of $c^{i}$)}.$ (83)
Consider a linear classifier with weight vector $\mathbf{w}$ applied to $\mathbf{f}_{x}(\mathbf{x})$:
$\text{logit}(\mathbf{x})=\mathbf{w}\cdot\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{w}\tilde{\mathbf{A}}^{i}\begin{bmatrix}p(c^{i}=0\mid\mathbf{x})\\ 1-p(c^{i}=0\mid\mathbf{x})\end{bmatrix}+\text{const}.$ (84)
To align this with the (unnormalized) output required for cross-entropy classification, we demand:
$\text{logit}(\mathbf{x})\approx s\cdot p(c^{i}=0\mid\mathbf{x})=s\cdot(1-p(c^{i}=1\mid\mathbf{x})),$ (85)
which implies:
$s_{0}-s_{1}\approx s,$ (86)
where $[s_{0},s_{1}]=\mathbf{w}\tilde{\mathbf{A}}^{i}$.
Ignoring the constant, we complete the proof. ∎
Appendix G Experimental Details Supporting Theoretical Results
G.1 Simulation Details
@@
G.2 Experimental Details on LLMs
@@
Generating such counterfactual pairs, however, presents significant challenges. First, the inherent complexity and nuances of natural language make it difficult to create pairs that differ in precisely the intended contexts while leaving other aspects unchanged. Second, as highlighted in prior works [59, 35], constructing such counterfactual sentences is a highly non-trivial task, even for human annotators, due to the intricacies of semantics and the need for precise control over contextual variations.
@@
To overcome these challenges, we adopt an approach based on existing resources. Specifically, we utilize the 27 counterfactual pairs introduced in [59], which provide a structured and well-curated set of counterfactual pairs. These concepts encompass a wide range of semantic and morphological transformations, as detailed in Table 1. By leveraging this established dataset, we ensure consistency with previous research while facilitating a robust and meaningful evaluation of Corollary 4.3.
@@
We first use these 27 counterfactual pairs to construct a $\mathbf{A}_{s}$ with size $27\times dim$ by using the differences in the representations of these 27 counterfactual pairs, where $dim$ corresponds to the feature dimension of the used LLM, such as 4096 for the Llama-27B model. To construct the corresponding matrix $\mathbf{W}_{s}$, we train a linear classifier using these 27 counterfactual pairs. Specifically, we use the representations of the counterfactual pairs as input and the corresponding values of the latent variables as output. As a result, the corresponding linear weights for the 27 counterfactual pairs are be used to create $\mathbf{W}_{s}$. In our experiments, we employ the LogisticRegression classifier from the scikit-learn library.
@@
Note that, as mentioned in Corollary 4.3, there is a scale $s$ in $\mathbf{A}_{s}\mathbf{W}_{s}\approx s\mathbf{I}$. Therefore, we normalize both $\mathbf{A}_{s}$ and $\mathbf{W}_{s}$, before computing $\mathbf{A}_{s}\times\mathbf{W}_{s}$. The scaling indeterminacy of latent variables is an inherent property of the latent space. However, in most cases, e.g., classification, the exact scaling factor is not particularly significant. For example, concepts are often represented as directions rather than magnitudes.
@@
Table 1: Concept names, one example of the counterfactual pairs, and the number of used pairs, taken from [59].
# Concept Example Word Pair Counts
1 verb $\Rightarrow$ 3pSg (accept, accepts) 50
2 verb $\Rightarrow$ Ving (add, adding) 50
3 verb $\Rightarrow$ Ved (accept, accepted) 50
4 Ving $\Rightarrow$ 3pSg (adding, adds) 50
5 Ving $\Rightarrow$ Ved (adding, added) 50
6 3pSg $\Rightarrow$ Ved (adds, added) 50
7 verb $\Rightarrow$ V + able (accept, acceptable) 50
8 verb $\Rightarrow$ V + er (begin, beginner) 50
9 verb $\Rightarrow$ V + tion (compile, compilation) 50
10 verb $\Rightarrow$ V + ment (agree, agreement) 50
11 adj $\Rightarrow$ un + adj (able, unable) 50
12 adj $\Rightarrow$ adj + ly (according, accordingly) 50
13 small $\Rightarrow$ big (brief, long) 25
14 thing $\Rightarrow$ color (ant, black) 50
15 thing $\Rightarrow$ part (bus, seats) 50
16 country $\Rightarrow$ capital (Austria, Vienna) 158
@@ 5 Ving $\Rightarrow$ Ved (adding, added) 50
18 male $\Rightarrow$ female (actor, actress) 52
19 lower $\Rightarrow$ upper (always, Always) 73
20 noun $\Rightarrow$ plural (album, albums) 100
21 adj $\Rightarrow$ comparative (bad, worse) 87
22 adj $\Rightarrow$ superlative (bad, worst) 87
23 frequent $\Rightarrow$ infrequent (bad, terrible) 86
24 English $\Rightarrow$ French (April, avril) 116
25 French $\Rightarrow$ German (ami, Freund) 128
26 French $\Rightarrow$ Spanish (annee, año) 180
27 German $\Rightarrow$ Spanish (Arbeit, trabajo) 228
Appendix H Experiment on Sparse Autoencoders
H.1 Implementation of the Proposed Structured SAE
The proposed structured SAE employs two regularization terms: a structured regularization to model the dependence among latent concepts, and a sparsity regularization based on the assumption that latent concepts may be sparsely activated. We implement it as follows:
$\mathbf{S}$ $=\text{ReLU}(\mathbf{w}_{s}(\mathbf{f}_{\mathbf{x}}(\mathbf{x})-\mathbf{b}_{d})+\mathbf{b}_{s}),$ (87)
$\mathbf{R}$ $=\text{ReLU}(\mathbf{w}_{l}(\mathbf{f}_{\mathbf{x}}(\mathbf{x})-\mathbf{b}_{d})+\mathbf{b}_{l}),$ (88)
$\mathbf{z}$ $=\mathbf{S}+\mathbf{R},$ (89)
$\mathbf{\bar{f}}(\mathbf{x})$ $=\mathbf{w}_{d}\mathbf{z}+\mathbf{b}_{d}.$ (90)
Here:
  • $\mathbf{S}$ denotes the sparse representations from the sparse encoder (with parameters $\mathbf{w}_{s},\mathbf{b}_{s}$);
  • $\mathbf{R}$ denotes the structured representation from the structured encoder (with parameters $\mathbf{w}_{l},\mathbf{b}_{l}$);
  • $\mathbf{z}$ is the combined representations used for reconstruction;
  • $\mathbf{\bar{f}}(\mathbf{x})$ denote the reconstruction of $\mathbf{f}(\mathbf{x})$.
The loss function for training is:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{train}}}\left[\|\mathbf{f}(\mathbf{x})-\mathbf{\bar{f}}(\mathbf{x})\|_{2}^{2}+\lambda_{t}\left(\|\mathbf{S}\|_{p_{t}}^{p_{t}}+\gamma\|\mathbf{R}\|_{\text{nuc}}\right)\right],$ (91)
where:
  • $\lambda_{t}$ is the dynamically adjusted sparsity coefficient at step $t$, following $p$-annealing SAE [38];
  • $\gamma$ is a hyperparameter that balances the sparsity penalty and the low-rank regularization;
  • $\|\mathbf{S}\|_{p_{t}}^{p_{t}}=\sum_{i}|s_{i}|^{p_{t}}$ is the adaptive $\ell_{p_{t}}$ norm promoting sparsity, following $p$-annealing SAE [38];
  • $\|\mathbf{R}\|_{\text{nuc}}$ is the nuclear norm, used to encourage low-rank structure.
We estimate the nuclear norm using the top-$k$ singular values:
$\|\mathbf{R}\|_{\text{nuc}}\approx\sum_{i=1}^{k}\sigma_{i},$ (92)
where $\{\sigma_{i}\}_{i=1}^{k}$ are the largest $k$ singular values, obtained via low-rank SVD (e.g., PyTorch’s svd_lowrank). The value of $k$ is a tunable parameter (e.g., $k=64$) that trades off approximation accuracy and computational cost.
H.2 Details of Evaluation Metric
Obtaining $p(c^{i}=1|\mathbf{x})$ from Supervised Linear Classification.
For evaluation, we first use 27 counterfactual pairs from [59], also see Table 1, to train a linear classification using LogisticRegression classifier from the scikit-learn library, to obtain logits, i.e., unnormalized $p(c^{i}=1|\mathbf{x})$, for each of the 27 pairs. As a result, for each concept in these 27 concepts, we can obtain the corresponding logit. Stacking these 27 logits yields the logit vector
$\mathbf{u}=\bigl{(}u_{1},u_{2},\dots,u_{27}).$ (93)
Extracting $z_{i}$ from trained SAEs.
We use the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ of the same 27 counterfactual pairs from [59] as input to a trained SAE, extracting the corresponding latent features $\mathbf{z}$. Let $\tilde{\mathbf{z}}$ denote the element-wise exponentiation of $\mathbf{z}$, i.e., $\tilde{\mathbf{z}}=\exp(\mathbf{z})$. This yields a feature matrix of size $27\times D$, where $D$ is the dimensionality of $\mathbf{z}$:
$\tilde{\mathbf{z}}=\bigl{(}\tilde{\mathbf{z}}_{1},\tilde{\mathbf{z}}_{2},\dots,\tilde{\mathbf{z}}_{27})^{T}.$ (94)
Correlation Matrix and Assignment.
For the logit vector $\mathbf{u}$, and the feature matrix $\tilde{\mathbf{z}}$, we compute the Pearson correlation
$\mathbf{R}_{d}=\operatorname{corr}\!\bigl{(}\mathbf{u},\,\tilde{\mathbf{z}}_{:,d}\bigr{)},\qquad d=1,\dots,D$ (95)
where $\tilde{\mathbf{z}}_{:,d}$ denotes the $d$-th column of $\tilde{\mathbf{z}}$.
Note that the estimated features $\mathbf{z}$ from the SAE are subject to permutation indeterminacy. To address this, we apply the Hungarian algorithm to solve the assignment problem on $\mathbf{R}_{d}$. This yields the optimal assignment for each concept, allowing us to compute the assigned Pearson correlation. We report the mean Pearson correlation across the 27 concepts, as well as the percentage of concepts whose Pearson correlation exceeds a threshold of 0.5.
H.3 Experiments and Results
We train four SAE variants—top-$k$ SAE [26], batch‑top‑$k$ SAE [17], $p$-annealing SAE [38], and our proposed structured SAE. Each SAE is trained on activations from the final hidden layer ($n=512$) of the 70 million-parameter Pythia language model [8], which was pre‑trained on The Pile corpus [25].
Experimental setup.
To obtain training data for the SAEs, we run the pretrained Pythia-70M model over a duplicated copy of The Pile and cache the representations of its last hidden layer. All SAE variants use the same feature dimension ($D=32{,}768$) and are trained for 20 000 optimization steps with a batch size of 10 000. We employ the Adam optimizer with an initial learning rate of $1\times 10^{-4}$ and linearly warm up the learning rate during the first 200 steps. For the top‑$k$ and batch-top-$k$ SAEs, we set $k=32$. For $p$-annealing SAEs, we apply a sparsity warm‑up of 400 steps and an initial sparsity penalty coefficient $\lambda_{s}=0.1$. The $p$-annealing‑LoRa SAE uses the same $p$-annealing settings and additionally applies a low-rank scaling factor $\gamma=0.1$.
Compute resources.
All experiments are conducted on a server equipped with four NVIDIA A100 GPUs (40 GB each) running CUDA 12.2 and driver version 535.161.07. The host system features an AMD EPYC 7313 (16 cores) CPU and 503 GB of RAM. Caching the last-layer activations of The Pile dataset requires 51.5 GPU-hours. Training the top-$k$ SAE takes 2 hours, while each of the other SAE variants requires 3 hours.
Figure 5: Classification accuracy of logistic probes across various concepts. Each bar represents the performance for a given concept.
Feature reconstruction.
To quantify how well each sparse auto‑encoder recovers the original hidden representations, we follow prior work and report the mean‑squared error (MSE) between the SAE‑reconstructed representations and the ground‑truth representations for the 27 evaluation word–pairs. Figure 6 plots the validation MSE over training steps for all four variants. All methods reduce error monotonically, but the two annealing strategies clearly dominate the fixed‑$k$ baselines. Our $p$‑annealing–LoRa SAE converges fastest and attains the lowest final MSE, followed by the plain $p$‑annealing SAE. Both top‑$k$ and Batch‑top‑$k$ remain larger the annealing approaches throughout training.
Figure 6: Validation reconstruction loss (lower is better) as training progresses.
Pearson Correlation Coefficient.
We report Pearson Correlation Coefficient (PCC) at the 20 000‑step checkpoint. As summarized in Table 2, the two $p$‑annealing variants markedly outperform the fixed‑$k$ baselines. Across the 27 concepts, the proposed structured SAE achieves the highest mean PCC (0.674) and covers 96.4% of concepts above the 0.5 threshold, followed closely by plain $p$‑annealing (0.638 / 89.3%). In contrast, top‑$k$ and batch‑top‑$k$ attain average PCCs of only 0.365 and 0.331, with fewer than 11% and 8% of concepts exceeding the 0.5 bar. The improvements are consistent: for 25 of the 27 concepts the strongest correlation is found in one of the annealing dictionaries, confirming that dynamic sparsity and the low‑rank adaptation yield features that align more cleanly with human‑interpretable concepts.
Table 2: PCC results at training step 20,000. Each row reports the PCC for a single concept, followed by the overall average and coverage@0.5 across the 27 concepts. Bold values denote the best‑performing method per row.
Paired‑concept correlation (PCC) across different SAEs ($\uparrow$)
Concepts top-$k$ batch-top-$k$ $p$-annealing Our structured SAE
3psg_ved 0.432 0.332 0.680 0.691
adj_adj_+_ly 0.379 0.396 0.794 0.833
adj_comparative 0.425 0.334 0.592 0.727
adj_superlative 0.321 0.302 0.772 0.746
adj_un_+_adj 0.534 0.472 0.741 0.732
average 0.365 0.331 0.638 0.674
country_capital 0.205 0.165 0.495 0.540
english_french 0.291 0.265 0.666 0.649
french_german 0.264 0.226 0.550 0.607
french_spanish 0.389 0.264 0.560 0.561
frequent_infrequent 0.262 0.255 0.472 0.554
german_spanish 0.250 0.251 0.667 0.701
lower_upper 0.204 0.293 0.531 0.543
male_female 0.413 0.312 0.532 0.544
noun_plural 0.357 0.346 0.636 0.650
pronoun_possessive 0.725 0.779 0.998 0.991
small_big 0.327 0.225 0.506 0.547
thing_color 0.588 0.574 0.895 0.880
thing_part 0.240 0.207 0.400 0.453
verb_3psg 0.312 0.268 0.622 0.617
verb_v_+_able 0.458 0.361 0.763 0.806
verb_v_+_er 0.318 0.363 0.609 0.666
verb_v_+_ment 0.407 0.288 0.623 0.674
verb_v_+_tion 0.429 0.357 0.653 0.706
verb_ved 0.371 0.333 0.573 0.656
verb_ving 0.369 0.485 0.612 0.728
ving_3psg 0.286 0.233 0.646 0.711
ving_ved 0.304 0.245 0.627 0.682
Average 0.365 0.331 0.638 0.674
Coverage@0.5 10.7% 7.1% 89.3% 96.4%
Ablation on the low‑rank coefficient.
We vary the low‑rank term scaling factor $\gamma\in\{0,10^{-3},10^{-2},10^{-1},1\}$ while holding every other hyper‑parameter fixed. The case $\gamma=0$ reduces to the plain $p$-annealing SAE baseline, whereas larger $\gamma$ values inject an increasingly strong low‑rank update. At each checkpoint we record three diagnostics: (i) reconstruction error (MSE), (ii) mean PCC, i.e., $\text{PCC}_{\text{avg}}$, and (iii) The percentage of concepts whose Pearson correlation exceeds a threshold of 0.5, i.e., coverage@0.5. Figure 7 overlays the trajectories of these metrics across training steps. A moderate adapter strength ($\alpha\approx 10^{-1}$) consistently yields the lowest MSE and the highest interpretability scores, while too large or too small values degrade at least one objective, underscoring the importance of tuning $\alpha$.
[…]
Appendix I Further Discussion: Observations on LLMs and World Models
I.1 LLMs Mimic the Human World Model, Not the World Itself